This one gets me because I think I'm close but cannot get a perfect match on the numbers.
I bought 4 bags of chips and 1 bag of peanuts for $1.75. Then my sister bough 2 bags of chips and 4 bags of peanuts for $2.00. Find the cost of a bage of chips and the cost for a bag of peanuts.
3x+1y=1.75
2x+4y=2.00
so,
3x+1y=1.75
-4x-8y=-400
-7y=-2.25
y=0.32
3x+1(0.32)=1.75
3x+0.32=1.75
3x=1.43
x=0.47

Answer by Paul(988)

Let the cost of bag of chips be x
Let the cost of bag of peanuts be y

4x+y=1.75----> y=1.75-4x (subsitution)

2x+4y=2
Subsitute)
2x+4(1.75-4x)=2
2x-16x=2-7
-14x=-5
x=0.36 (rounded)

y=1.75-4(0.36)
y=0.31

Hence, a cost of a bag of chips is $0.36 and a cost of a bag of peanuts is $0.31.
Paul.